# Leveraging Joint Predictive Embedding and Bayesian Inference in Graph Self Supervised Learning

### v2

@@ IV-B5 Study on Performance against Test Time Node Feature Distortion
We evaluate the performance of our method on node classification by augmenting the test splits. We sample $p$ percentage test nodes randomly and replace their node features with points sampled from a standard Gaussian distribution. We evaluate the performance of the distorted test nodes on Amazon Computers, WikiCS, Amazon Photos and Coauthor CS by varying $p$ between 10% and 40% and report their performance differences in table VII. In this study, we only use WikiCS, Co-author and Amazon networks since Planetoid datasets tend to be unstable on evaluation as indicated in [21]. Despite no further fine-tuning on distorted node features, the average performance drop on Amazon Photos and Computers is only 2.96% and 4.01% respectively. This indicates the ability of the joint framework to generalize on noise-augmented views as well. According to our hypothesis, we believe this is due to the nature of joint predictive embedding learning where a higher variation in context and target embeddings is required for better generalization.

### v4

@@ IV-B5 Study on Performance against Test Time Node Feature Distortion
We evaluate the performance of our method on node classification by augmenting the test splits. We sample $p$ percentage test nodes randomly and replace their node features with points sampled from a standard Gaussian distribution. We evaluate the performance of the distorted test nodes on Amazon Computers, WikiCS, Amazon Photos and Coauthor CS by varying $p$ between 10% and 40% and report their performance differences in table VII. In this study, we only use WikiCS, Co-author and Amazon networks since Planetoid datasets tend to be unstable on evaluation as indicated in (shchur2018pitfalls, ). Despite no further fine-tuning on distorted node features, the average performance drop on Amazon Photos and Computers is only 2.96% and 4.01% respectively. This indicates the ability of the joint framework to generalize on noise-augmented views as well. According to our hypothesis, we believe this is due to the nature of joint predictive embedding learning where a higher variation in context and target embeddings is required for better generalization.